OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 2x + 3y = 3 3x – 2y = 11 What is the solution to the system? (–3, 3) (–1, –7) (1, –4) (3, –1)

OpenStudy (anonymous):

2/ 2x + 3y = 3 3/ 3x – 2y = 11 4x+6y=6 9x-6y=33 +-------- 13x=39 x=3

OpenStudy (anonymous):

Graph of the system:

OpenStudy (anonymous):

Performed with the free online graphing tool at: http://www.meta-calculator.com/online/

OpenStudy (anonymous):

The first equation is the blue line, the second is the red line.

OpenStudy (anonymous):

I still dont understand how to do the problem though

OpenStudy (anonymous):

The problem looks like you are to get the solution graphically. You can always confirm the answer algebraically, too. Is it that you don't know how to use the graphing tool or are you saying that you don't understnad the lines drawn?

OpenStudy (anonymous):

honestly i dont understand any of it at all